If a 75 m wire is cut into three pieces having proportions 2:4:6 then the length ,in meters ,of the smallest piece is
A 61/4 b 121/2 c 131/3 d 162/3 e none of these

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

The total number of pieces is 12
1/12 of 75 = 75/12 meters
2parts = 150 /12----12.5 meters
4 parts = 300/12-----25 meters
6 parts = 450/12-------37.5 meters

Answer by richwmiller(17219)

There aren't 12 pieces but only three
there are 12 parts to the proportion
2x+4x+6x=12x
one piece has 2 of those 12 parts
one piece has 4 parts
one piece has 6 parts
one part=6.25
2 parts=12.5
4 parts 25
6 parts 37.5
the smallest piece is 12.5
Answer e) none of these
61/4 is 15.25
121/2=60.5
131/3=43 2/3
162/3=54
the smallest piece is 12.5
multiplied by 2 is 25
multiplied by 3 is 37.5
multiplied by 4 is 50
none of these appear as a numerator